Henderson has 5,000 defective television on hand, which cost $400,000 to manufacture. Henderson can either sell these defective televisions as scrap for $65 per unit, or spend an additional $120,000 on repairs and then sell the televisions for $130 per unit.

Should Henderson repair the defective TV's to sell them as scrap?

You must show your supporting computations
Answer #1

Yes, Henderson should repair the defective TV's to sell them as scrap.

Manufacturing cost per unit of defective television (without repair) = 400000/5000 = $80

Scrap value per unit of defective television (without repair) = $65

Loss per unit on sell (without repair) =manufacturing cost –sell price = 80-65 = $15

Manufacturing cost per unit of defective television (with repair)= (400000+120000)/5000 = $104

Scrap value per unit of defective television (with repair) = $130

Profit per unit on sell (with repair) = sell price - manufacturing cost = 130-104 = $26

Total profit = profit per unit * number of units = 26*5000 = $130000